OpenStudy (anonymous):

Which ordered pairs in the form (x, y) are solutions to the equation 4x – 5y = 24? Choose all answers that are correct. A. (−9, −12) B. (1, −4) C. (4, 8) D. (6, 0)

OpenStudy (fateal):

like this?

OpenStudy (justuu):

Hi, Since parallel lines have the same slope, first find the slope of 4x + 5y = 24. 4x + 5y = 24 5y = -4x + 24 y = -4/5x + 24/5 In y = mx + b form, the slope is -4/5. To find the equation of the parallel line through (0,12), 12 is its y intercept or value of "b" in y = mx + b. So, the equation is: y = -4/5x + 12 or y = -.8x + 12 To find the x intercept of this line, let y = 0 and then solve for x. y = -4/5x + 12 0 = -4/5x + 12 4/5x = 12 4x = 60 x = 15 So when y = 0, then x = 15 gives the x intercept as (15,0). <==ANSWER I hope that helps!! :-)

OpenStudy (justuu):

MAYBE THIS ONE IS BETTER Solution for 4x-5y=24 equation: Simplifying 4x + -5y = 24 Solving 4x + -5y = 24 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '5y' to each side of the equation. 4x + -5y + 5y = 24 + 5y Combine like terms: -5y + 5y = 0 4x + 0 = 24 + 5y 4x = 24 + 5y Divide each side by '4'. x = 6 + 1.25y Simplifying x = 6 + 1.25y

OpenStudy (anonymous):

in other words it is a, b, d